# The Future of Large Language Model Pre-training is Federated

### v2

@@ Abstract
Generative pre-trained large language models (LLMs) have demonstrated impressive performance over a wide range of tasks, thanks to the unprecedented amount of data they have been trained on. As established scaling laws indicate, LLMs’ future performance improvement depends on the amount of computing and data sources they can leverage for pre-training. Federated learning (FL) has the potential to unleash the majority of the planet’s data and computational resources, which are underutilized by the data-center-focused training methodology of current LLM practice. Our work presents a robust, flexible, reproducible FL approach that enables large-scale collaboration across institutions to train LLMs. We propose a scalable deployment system called Photon to enable the investigation and development of this new training paradigm for LLM pre-training. We show that Photon can be used by organizations interested in collaborating with their private data sources and computational resources for pre-training LLMs with billions of parameters. This paradigm would mobilize more computational and data resources while matching or potentially exceeding centralized performance. We further show the effectiveness of the federated training scales with model size and present our approach for training a billion-scale federated LLM using limited resources. Finally, we show that LLM training is highly resilient to the classical challenges of federated statistical and hardware heterogeneity. Furthermore, we show that convergence is robust to partial participation, opening the avenue for compute-efficient collaborative training. Photon will help data-rich actors to become the protagonists of LLMs pre-training instead of leaving the stage to compute-rich actors alone.
@@ 1 Introduction
We are the first [1] to fully utilize a practical and production-ready geographically distributed federation of heterogeneous devices and data sources for reproducible generative pre-training of a billion-scale LLM. This work presents Photon, a complete system for pre-training federated LLMs in a collaborative, reproducible, and scalable fashion that can address the main challenges for a flexible execution in arbitrary FL cross-silo settings. We built it on the open-source FL framework Flower [23]. Furthermore, we make this available as a training recipe with a fully transparent experimental configuration accompanying our results. We emphasize that federated generative pre-training can be done with affordable hardware configurations by interlinking single nodes containing $1$-$8$ GPUs from standard cloud providers —rather than renting multiple entire data centers concurrently. We also argue that federated systems can and must evolve beyond networks of clients holding both compute and data in a world of increasing hardware costs and competing privacy interests. To achieve this, Photon allows data providers (sources) to come together with compute providers in private partnerships amongst trusted parties without leaking any such data to other participants. We demonstrate that pre-training LLMs can be democratized through FL technologies, including data-rich actors independent of their connectivity and computing resources.
@@
    Federated LLM training offers competitive performance with centralized training but with far less communication overhead.
@@ 6.3 Experimental Design
First, we wish to isolate the impact of the optimization procedure itself by using a standard benchmark dataset partitioned across clients in an IID fashion. Since parameter averaging is known to inject noise into the optimization procedure [18], it may cause a change in the dynamics of LLM training. Thus, we want to observe whether it can substitute distributed data-parallel SGD for situations where data is mainly similar from client to client (e.g., in the same language and containing the same mix of genres). We use a version of C4 [80] randomly split into randomly constructed shards for our homogeneous data sources experiments.
@@
Second, we want to observe whether injecting federated statistical heterogeneity into the system harms convergence. We are particularly concerned with approximating a realistic cross-silo form of data heterogeneity similar to those encountered by institutions like publishers, which may specialize in different text genres. For this purpose, we choose to use The Pile partitioned into subsets representing general knowledge (Wikipedia, English only), general scientific publications (ArXiv), long-form books  (Project Gutenberg, alias PG-19), technical news/discussions (HackerNews), medical publications (PubMed Central), law (FreeLaw), philosophical publications (PhilPapers), and Q/A websites (StackExchange). This coarse mapping allows us to directly determine the efficacy of federated training in reconciling different styles and sources of knowledge.
Model Scaling:
Third, we want to show that federated optimization can train a large model in a cost-effective distributed fashion without access to the sprawling data centers traditionally tasked with such training. Given the prohibitive nature of such training, we initially focus on the IID $C4$ partition due to its usefulness as a generally agreed-upon benchmark in the community.
@@
Table 1: Architecture details and local training parameters for our $75$M, $125$M, $350$M and $1.3$B models. We report the number of transformer blocks, hidden model dimension $d$, number of attention heads, the linear layer expansion ratio, and Adam’s parameters ($\beta_{1}$ and $\beta_{2}$). We also report the vocabulary size of the tokenizer we used [101].
Model Size #Blocks $\boldsymbol{d}$Heads Exp. Ratio $\mathbf{(\boldsymbol{\beta_{1}},\leavevmode\nobreak\ \boldsymbol{\beta_{2}})}$ $\boldsymbol{|\textbf{Vocab}|}$
@@ 6.4 Training Setup
All experiments use $500$ local steps per round, executed by all the clients in the federation. We distinguish between parallel steps and sequential steps since the model parameters are averaged after each round, and each of the $500$ local steps in a round is done in parallel, as in standard data-parallel training [34]. Our setting respects the conditions of standard cross-silo [50] FL in which the orchestrator samples the entire batch of equally capable clients in every round. Our experiments used diverse hardware resources whose specifications allow for reasonably fast execution; as an example, we used a combination of heterogeneous servers equipped with NVIDIA A40, A100, and H100 GPUs. Due to FL’s lax synchronization requirements, these heterogeneous hardware accelerators were able to collaborate despite being located in different countries.
As shown in Table 1, we trained models ranging in size from $75$ million parameters to $1.3$ billion for the causal language modeling task. We used the tokenizer presented in [101] with a vocabulary size of 50 3685036850\,36850 368. The local optimizer the clients use in our experiments is AdamW [102], while the server optimizer is FedMom [47]. The hyperparameters we used are reported in Table 2.
@@
Model Size #Rounds $\boldsymbol{\eta_{s}}$ $\boldsymbol{\mu_{s}}$ $\boldsymbol{\alpha}$ $\boldsymbol{\eta_{max}}$ $\boldsymbol{T}$ Batch Size
$\mathbf{75}$M $40$ $0.2$ $0.9$ $10^{-6}$ $4\times 10^{-4}$ 88 0008800088\,00088 000 $256$
$\mathbf{125}$M $25$ $0.2$ $0.9$ $10^{-5}$ $3\times 10^{-4}$ 15 0001500015\,00015 000 $256$
$\mathbf{350}$M $40$ $0.2$ $0.9$ $10^{-1}$ $3\times 10^{-4}$ 13 4001340013\,40013 400 $256$
$\mathbf{1.3}$B $16$ $0.2$ $0.9$ $10^{-1}$ $2\times 10^{-4}$ 24 8002480024\,80024 800 $512$
@@ 7 Evaluation
We have experimentally validated the efficacy of our federated training recipe at a series of increasing model scales while investigating impactful choices for both the federated and local optimization. Thus, the results presented in this section indicate the viability of federated optimization for LLM pre-training and validate its ability to reach beyond the data and hardware constraints of centralized approaches.
@@
Federated Optimization is Effective and Stable across Scales:
Our initial experiments using the IID partition of the C4 dataset reveal that federated optimization can reach a performance comparable to centralized training. As can be seen in Fig. 3, federated models perform comparatively well to centrally trained ones or even better at sufficiently large model sizes. Furthermore, while their convergence curves may initially look more unstable, the federated aggregation procedure quickly resolves this once the model settles into an appropriate compromise for the clients. Most importantly, FL effectively trains the $\mathbf{1.3}$B parameters model. The gap between centralized and federated performance becomes increasingly small as the model size increases, ranging between $10$ and $2$ points on the perplexity axis.
@@
Federated Optimization is Robust to Heterogeneity:
@@
Larger Models Improve Consensus:
@@
Federated Optimization is Robust to Partial Participation:
@@
Interplay of Client and Server Models:
@@
Federated Optimization Aligns Client Gradients:

### v3

@@ Abstract
Generative pre-trained large language models (LLMs) have demonstrated impressive performance over a wide range of tasks, thanks to the unprecedented amount of data they have been trained on. As established scaling laws indicate, LLMs’ future performance improvement depends on the amount of computing and data sources they can leverage for pre-training. Federated learning (FL) has the potential to unleash the majority of the planet’s data and computational resources, which are underutilized by the data-center-focused training methodology of current LLM practice. Our work presents a robust, flexible, reproducible FL approach that enables large-scale collaboration across institutions to train LLMs. We propose a scalable deployment system called Photon to enable the investigation and development of this new training paradigm for LLM pre-training. We show that Photon can be used by organizations interested in collaborating with their private data sources and computational resources for pre-training LLMs with billions of parameters. This paradigm would mobilize more computational and data resources while matching or potentially exceeding centralized performance. We further show the effectiveness of the federated training scales with model size and present our approach for training billion-scale federated LLMs using limited resources. Thus far, we have used Photon to train LLM models to the size of $\mathbf{7}$B parameters and anticipate larger models being completed in the near future. Finally, we show that LLM training is highly resilient to the classical challenges of federated statistical and hardware heterogeneity. Furthermore, we show that convergence is robust to partial participation, opening the avenue for compute-efficient collaborative training. Photon will help data-rich actors to become the protagonists of LLMs pre-training instead of leaving the stage to compute-rich actors alone.
@@ 1 Introduction
We are the first [1] to fully utilize a practical and production-ready geographically distributed federation of heterogeneous devices and data sources for reproducible generative pre-training of billion-scale LLMs. This work presents Photon, a complete system for pre-training federated LLMs in a collaborative, reproducible, and scalable fashion that can address the main challenges for a flexible execution in arbitrary FL cross-silo settings. We built it on the open-source FL framework Flower [23]. Furthermore, we make this available as a training recipe with a fully transparent experimental configuration accompanying our results. We emphasize that federated generative pre-training can be done with affordable hardware configurations by interlinking single nodes containing $1$-$8$ GPUs from standard cloud providers —rather than renting multiple entire data centers concurrently. We also argue that federated systems can and must evolve beyond networks of clients holding both compute and data in a world of increasing hardware costs and competing privacy interests. To achieve this, Photon allows data providers (sources) to come together with compute providers in private partnerships amongst trusted parties without leaking any such data to other participants. We demonstrate that pre-training LLMs can be democratized through FL technologies, including data-rich actors independent of their connectivity and computing resources.
@@
    Federated LLM training offers competitive performance with centralized training but with far less communication overhead in a broad range of model sizes up to $\mathbf{7}$B parameters.
@@ 6.3 Experimental Design
We wish to isolate the impact of the optimization procedure itself by using a standard benchmark dataset partitioned across clients in an IID fashion. Since parameter averaging is known to inject noise into the optimization procedure [18], it may cause a change in the dynamics of LLM training. Thus, we want to observe whether it can substitute distributed data-parallel SGD for situations where data is mainly similar from client to client (e.g., in the same language and containing the same mix of genres). We use a version of C4 [80] randomly split into randomly constructed shards for our homogeneous data sources experiments.
@@
We want to observe whether injecting federated statistical heterogeneity into the system harms convergence. We are particularly concerned with approximating a realistic cross-silo form of data heterogeneity similar to those encountered by institutions like publishers, which may specialize in different text genres. For this purpose, we choose to use The Pile partitioned into subsets representing general knowledge (Wikipedia, English only), general scientific publications (ArXiv), long-form books  (Project Gutenberg, alias PG-19), technical news/discussions (HackerNews), medical publications (PubMed Central), law (FreeLaw), philosophical publications (PhilPapers), and Q/A websites (StackExchange). This coarse mapping allows us to directly determine the efficacy of federated training in reconciling different styles and sources of knowledge.
@@
6.4 Model Scaling
We want to show that federated optimization can train a large model without access to the sprawling data centers traditionally tasked with such training. Given the prohibitive nature of such training, we initially focus on the IID $C4$ partition, as it represents a generally agreed-upon benchmark in the community.
We inspire from the scaling laws of Hoffmann et al. [6] to reason about what could be an equipollent discussion for federated training. In particular, we are interested in making consistent choices regarding the number of tokens the models train on. In centralized settings, the estimate for an optimal ratio, $\frac{\mathrm{dim}(\mathcal{D})}{\mathrm{dim}(\Theta)}$, is 20 tokens/parameter, with $\mathrm{dim}(\mathcal{D})$ being the number of tokens of the dataset $\mathcal{D}$ and $\mathrm{dim}(\Theta)$ being the number of trainable parameters for the model $\Theta$. However, we must consider that Hoffmann et al. [6] used a tokenizer with about $32$K tokens in its vocabulary, while we intend to use $50$K. This different tokenization impacts the size of the embedding layer only. If we assume that the scaling laws of Hoffmann et al. [6] will hold even with a larger tokenizer, we can scale down the size of the model we train to obtain their Hoffmann et al. [6] equivalents.
While the scaling laws proposed by Hoffmann et al. [6] provide a compute-optimal way to achieve a specific loss, they do not account for inference costs after pre-training. Sardana et al. [101], affiliated with MosaicML, proposes a longer pre-training phase to improve inference performance, which . Very recently, He et al. [102] investigated the particular case of LocalSGD in training LLMs. Despite that work being admittedly limited, we hope that future research in this direction will clarify optimal choices for training LLMs with this approach. Our estimates are presented in Table 1.
Table 1: Pre-training tokens and steps for a given model size $\mathrm{dim}(\theta)$ in parameters, the size in the parentheses represents the vocabulary-adjusted size to match Hoffmann et al. [6]. Where $\mathrm{dim}(\mathcal{D})|_{\Theta}$ is the compute-optimal number of tokens recommended by Hoffmann et al. [6], $\mathrm{dim}(\mathcal{D})_{\mathrm{MPT}}|_{\Theta}$ is the number recommended by MosaicML. We use $\mathrm{dim}(\mathcal{D})_{\mathrm{SEQ}|\theta}^{*}$ as the number of sequential tokens and $\mathrm{dim}(\mathcal{D})_{\mathrm{PAR}|\theta}^{*}$ as the number of parallel ones, which depends on the number of clients. Given the sequence length and batch size $l$, $B$, we report the equivalent number of steps $\mathcal{T}$.
$\mathrm{dim}(\Theta)$ $\mathrm{dim}(\mathcal{D})|_{\Theta}$ $\mathrm{dim}(\mathcal{D})_{\mathrm{MPT}}|_{\Theta}$ $\mathrm{dim}(\mathcal{D})_{\mathrm{SEQ}|\theta}^{*}$ $\mathrm{dim}(\mathcal{D})_{\mathrm{PAR}|\theta}^{*}$ $l$ $B$ $\mathcal{T}_{\mathrm{dim}(\mathcal{D})|\Theta}$ $\mathcal{T}_{\mathrm{dim}(\mathcal{D})_{\mathrm{MPT}}|_{\Theta}}$ $\mathcal{T}_{\mathrm{dim}(\mathcal{D})_{\mathrm{SEQ}|\theta}^{*}}$
$75\mathrm{M}\,(58.54\mathrm{M})$ $1.17\times 10^{9}$- $5.2\times 10^{9}$ $41.9\times 10^{9}$ $1024$ $256$ $4463$- $88000$
$125\mathrm{M}\,(110.89\mathrm{M})$ $2.22\times 10^{9}$ $2.5\times 10^{9}$ $6.6\times 10^{9}$ $52.4\times 10^{9}$ $2048$ $256$ $4235$ $4800$ $15000$
$350\mathrm{M}\,(331.19\mathrm{M})$ $6.62\times 10^{9}$ $8.0\times 10^{9}$ $10.5\times 10^{9}$ $83.9\times 10^{9}$ $2048$ $256$ $12627$ $15360$ $13500$
$1.3\mathrm{B}\,(1.26\mathrm{B})$ $25.2\times 10^{9}$ $26.0\times 10^{9}$ $7.35\times 10^{9}$ $58.8\times 10^{9}$ $2048$ $512$ $24033$ $24800$ $7000$
$3\mathrm{B}\,(2.96\mathrm{B})$ $59.2\times 10^{9}$ $54.0\times 10^{9}$ $13.1\times 10^{9}$ $52.4\times 10^{9}$ $2048$ $512$ $56458$ $51500$ $10500$
$7\mathrm{B}\,(6.92\mathrm{B})$ $138\times 10^{9}$ $134.0\times 10^{9}$ $22.0\times 10^{9}$ $88.1\times 10^{9}$ $2048$ $1024$ $65804$ $63900$ $10500$
Table 2: Architecture details and local training parameters for our $75$M, $125$M, $350$M, $1.3$B, $3$B, and $7$B models. We report the number of transformer blocks, hidden model dimension $d$, number of attention heads, the linear layer expansion ratio, and Adam’s parameters ($\beta_{1}$ and $\beta_{2}$). We also report the vocabulary size of the tokenizer we used [103] and the sequence length $l$.
Model #Blocks $\boldsymbol{d}$Heads Exp. Ratio $\mathbf{(\boldsymbol{\beta_{1}},\leavevmode\nobreak\ \boldsymbol{\beta_{2}})}$ $\boldsymbol{|\textbf{Vocab}|}$ $\boldsymbol{l}$
Size
[…]
@@ 6.5 Training Setup
All experiments use $500$ local steps per round, executed by all the clients in the federation.
Our setting respects the conditions of standard cross-silo [50] FL in which the orchestrator samples the entire batch of equally capable clients in every round. Our experiments used diverse hardware resources whose specifications allow for reasonably fast execution; as an example, we used a combination of heterogeneous servers equipped with NVIDIA A40, A100, and H100 GPUs. Due to FL’s lax synchronization requirements, these heterogeneous hardware accelerators were able to collaborate despite being located in different countries.
As shown in Table 2, we trained models ranging in size from $75$ million parameters to $7$ billion for the causal language modeling task. We used the tokenizer presented in [103] with a vocabulary size of 50 3685036850\,36850 368. The local optimizer the clients use in our experiments is AdamW [104], while the server optimizer is FedMom [47]. The hyperparameters we used are reported in Table 3.
@@
Model $\boldsymbol{\eta_{s}}$ $\boldsymbol{\mu_{s}}$ $\boldsymbol{\alpha}$ $\boldsymbol{\eta_{max}}$ $\boldsymbol{T}$ Batch
Size Size
$\mathbf{75}$M $0.7$ $0.9$ ${10^{-1},10^{-6}}$ $4\times 10^{-4}$ 88 0008800088\,00088 000 $256$
$\mathbf{125}$M $0.3,0.5,0.7$ $0.9$ ${10^{-1},10^{-5}}$$\{3.0,6.0\}\times 10^{-4}$ 15 0001500015\,00015 000 $256$
$\mathbf{350}$M $0.1$ $0.9$ $10^{-1}$ $3\times 10^{-4}$ 13 4001340013\,40013 400 $256$
$\mathbf{1.3}$B $0.7$ $0.9$ $10^{-1}$ $2\times 10^{-4}$ 24 8002480024\,80024 800 $512$
$\mathbf{3}$B $0.7$ $0.9$ $10^{-1}$ $1.6\times 10^{-4}$ 51 5005150051\,50051 500 $512$
$\mathbf{7}$B $0.7$ $0.9$ $10^{-1}$ $1.2\times 10^{-4}$ 63 9006390063\,90063 900 $1024$
Table 4: Hyperparameters for our federated experiments. $P$ represents the total number of clients per federations, $K$ the number of clients sampled per round, $D$ the dataset, $\tau$ the number of steps per round.
Model #Rounds $\boldsymbol{P}$ $\boldsymbol{K}$ $\boldsymbol{D}$ $\boldsymbol{\tau}$
Size
$\mathbf{75}$M $40$ ${8,64}$ ${8,4}$ C4 [80], The Pile [99] $500$
$\mathbf{125}$M ${10,25}$ ${8,64}$ ${8,4}$ C4 [80], The Pile [99] ${250,500}$
$\mathbf{350}$M $40$ $8$ $8$ C4 [80] $500$
$\mathbf{1.3}$B $14$ $8$ $8$ C4 [80] $500$
$\mathbf{3}$B $21$ $64$ $4$ C4 [80] $500$
$\mathbf{7}$B $21$ $64$ $4$ C4 [80] $500$
@@ 7 Evaluation
We have experimentally validated the efficacy of our federated training recipe at a series of increasing model scales while investigating impactful choices for both the federated and local optimization. Thus, the results presented in this section indicate the viability of federated optimization for LLM pre-training and validate its ability to reach beyond the data and hardware constraints of centralized approaches. We push our evaluation from relatively small sizes all the way to $7$B parameters in Section 7.7.
@@
7.1 Federated Optimization is Effective and Stable across Scales:
Our initial experiments using the IID partition of the C4 dataset reveal that federated optimization can reach a performance comparable to centralized training. As can be seen in Fig. 3, federated models perform comparatively well to centrally trained ones or even better at sufficiently large model sizes. Furthermore, while their convergence curves may initially look more unstable, the federated aggregation procedure quickly resolves this once the model settles into an appropriate compromise for the clients. Most importantly, FL effectively trains the $\mathbf{1.3}$B parameters model. The gap between centralized and federated performance becomes increasingly small as the model size increases, ranging between $10$ and $2$ points on the perplexity axis. In Fig. 9, we extend our experiments to show that this trend not only continues for $3$B and $7$B but accelerates.
@@ 7.1 Federated Optimization is Effective and Stable across Scales:
7.2 Federated Optimization is Robust to Heterogeneity:
@@
7.3 Larger Models Improve Consensus:
@@
7.4 Federated Optimization is Robust to Partial Participation:
@@
7.5 Interplay of Client and Server Models:
@@
7.6 Federated Optimization Aligns Client Gradients:
@@
7.7 Pushing the Boundary on Federated LLM Size
Having established that larger models reach consensus faster and that federated pre-training can achieve comparable perplexity to centralized pre-training, we now aim to examine how this behavior extrapolates to even larger sizes. To this end, we used Photon to pre-train $3$B and $7$B models using the same recipe.
As shown in Fig. 9, these larger sizes consistently outperform their centralized counterparts regarding client training perplexity and server validation perplexity. The gap in validation perplexity is larger than the one previously obtained for the $1.3$B, suggesting continued improvements in the effectiveness of federated optimization.
Figure 9: Comparison between the perplexity of the federated global model evaluated on the centralized validation set, the train perplexities of federated clients (averaged together), and the train and test perplexities for a centralized experiment. These metrics are reported for our $3$B (a) and $7$B (b) experiments. These large models reach a lower perplexity than their centralized counterparts and are more robust in the face of federated aggregation, with minimal perplexity spikes after the early rounds.
7.8 On the choice of the outer optimizer
Figure 10: Training cross-entropy comparison for (a) large batches and (b) small batch sizes. In both conditions, standard FedAvg outperforms SGD+N (SGD + server-side Nesterov momentum) and FedAvg-KeepOpt (FedAvg with local optimizer states saved between rounds). As can be seen in (c) and (b), the effect of server-side momentum and of keeping local optimizer states is to cause growth in the l2 norm of the model, causing the eventual divergence in cross-entropy.
Finally, it is worth considering several of the relevant choices that emerge from the interplay between local and federated optimization. The first of these is the choice of server-side and local optimizers. For the local optimizer, we concern ourselves with AdamW [104], assuming to preserve the most popular choice of the standard distributed approaches for our local training pipeline. For the server-side optimizer, we consider FedAvg [19] and server-side Nesterov momentum [47] as used in Douillard et al. [20]. A final choice worth considering when using a stateful local optimizer is whether to preserve optimizer states in between rounds.
The second choice we are concerned with is the local batch size. Since increasing the number of clients behaves similarly to increasing batch size [18], we consider two scenarios: (a) keeping the same batch size locally as the centralized pre-training recipe, (b) decreasing the local batch size such that the effective batch size, number of clients times the local batch size, matches that of centralized pre-training.
Our results, shown in Fig. 10, indicate that standard FedAvg provides the lowest final perplexity and is the most robust to divergence without keeping local optimizer states. While both server-side momentum and keeping optimizer states can provide initial speedups, they fail to approach the perplexity of centralized training. The same trends hold when using smaller batches locally at the expense of greater communication frequency.
Beyond performance benefits, not maintaining optimizer states also enables stateless clients, allowing, for example, partial participation without staleness, tolerating dropouts, and admitting new entrants into the federations. Furthermore, keeping optimizer states under data heterogeneity may prove cumbersome, as it increases client drift without synchronization. If optimizer states are synchronized, communication costs can grow threefold. Given these factors, we prefer and recommend stateless clients over the approach of previous methods [20].
7.9 Evaluation Harness
Table 5: In-context learning comparison between Photon models. Our biggest model wins $6$ out of $7$ comparisons in this group.
Name
ARC-Challenge
[108]
BigBench
QA Wikidata
[109]
HellaSwag
[110]
PIQA
[111]
Winogrande
[112]
ARC-Easy
[108]
BoolQ
[113]
Photon-7B $\mathbf{0.265}$ $\mathbf{0.447}$ $\mathbf{0.524}$ $\mathbf{0.729}$ 0.522 $\mathbf{0.508}$ 0.530
Photon-3B 0.247 0.360 0.455 0.705 0.512 0.461 0.591
Photon-1B 0.243 0.215 0.349 0.676 0.516 0.390 $\mathbf{0.612}$
Table 6: In-context learning comparison between Photon models. Our biggest model wins $5$ out of $6$ comparisons in this group.
Name
Openbook QA
[114]
Winograd
[115]
LAMBADA
(OpenAI) [116]
Bigbench [109]
Strategy QA
COPA
[117]
MMLU
[118]
Photon-7B $\mathbf{0.358}$ $\mathbf{0.681}$ $\mathbf{0.457}$ 0.466 $\mathbf{0.710}$ $\mathbf{0.263}$
Photon-3B 0.316 0.656 0.381 0.464 0.620 0.252
Photon-1B 0.274 0.604 0.298 0.470 0.630 0.248
To evaluate the downstream task performance of our models, we test across a series of in-context learning benchmarks. Our results, shown in Tables 5 and 6, demonstrate that the downstream performance of models trained with Photon scales as expected with model size, with our largest model winning $10$ out of $14$ reasons. This proves the downstream utility of Photon models even when using a pre-training dataset not optimized for downstream performance. We expect that as we continue to increase the model size and incorporate a broader and more qualitative data mixture, the downstream performance of Photon models will keep improving.